# A Survey of Text Watermarking in the Era of Large Language Models

### v2

@@ 4.3. Watermarking during Token Sampling
The previous section primarily focused on incorporating watermarks during the logits generation phase for Large Language Models. However, even after embedding watermarks, it is necessary to sample the next token from the watermarked logits. Consequently, the effectiveness of the watermark might be influenced by different sampling methods; for instance, beam search typically allows for a higher watermark intensity. In this section, we will introduce a technique of watermarking during token sampling, which does not alter the logits produced by the LLM. The primary advantage of this method is that the generated text is usually unbiased, thus minimally affecting text quality.
@@
To address this issue, Kuditipudi et al. (2023) proposed the use of a pseudo-random number sequence significantly longer than the text, randomly selecting a starting position from the sequence for each watermark insertion to introduce randomness. Additionally, during watermark detection, they incorporate a soft notion of edit distance (i.e., Levenshtein distance) into the computation of the alignment between text and the pseudo-random number sequence. This approach significantly enhances the robustness of the watermarking algorithm against watermark removal attacks. Currently, there is limited research on watermarking during token sampling, suggesting substantial potential for growth in this field. Moreover, the practical effectiveness and robustness of this method may require further validation through more experiments and real-world applications.
@@ 5. Evaluation Metrics for Text Watermarking
Figure 4. A more illustrative explanation of four distinct testing perspectives for text watermarking algorithms: success rate (section 5.1), text quality (section 5.2), robustness (section 5.3), and unforgeability (section 5.4).
In sections 3 and 4, we have provided a comprehensive introduction to the existing text watermarking methods. Meanwhile, for a text watermarking algorithm, it is crucial to conduct a comprehensive evaluation from various perspectives. In this section, we will introduce how to evaluate a watermarking algorithm from four perspectives: success rate, text quality, robustness and unforgeability. Among these, success rate (section 5.1) refers to the ability to correctly detect watermarked texts. Text quality (section 5.2) assesses whether the quality of watermarked text is degraded compared to non-watermarked text. Robustness (section 5.3) examines whether the watermarked text can still be detected after watermark removal attacks. Finally, unforgeability (section 5.4) evaluates the difficulty for third parties to forge the watermark. Meanwhile, we also provide a more illustrative depiction of the different evaluation methods in Figure 4.
@@ 5.1.1. Zero-bit Watermark
In zero-bit watermarking, the watermarking algorithm can only determine whether a text contains a watermark, without the ability to extract additional information from the watermarked text. Given that the detection of a zero-bit watermark inherently represents a binary classification problem — discerning whether a text is watermarked or not — the evaluation of its success rate typically employs classification metrics, such as accuracy and F1 score. In most works (Kirchenbauer et al., 2023b; Zhao et al., 2023a; Liu et al., 2023b, c), since the proportion of watermarked and non-watermarked texts in the test dataset is usually equal, the values of accuracy and F1 score tend to be quite similar. Despite this, the F1 score is often more frequently utilized, accompanied by additional metrics including false positive and false negative rate (Kirchenbauer et al., 2023b; Zhao et al., 2023a; Liu et al., 2023b). Here, false positives denote the erroneous classification of non-watermarked texts as watermarked, whereas false negatives refer to the incorrect classification of watermarked texts as non-watermarked. Generally, false positives are more important since misidentifying human-generated texts as watermarked can lead to more adverse consequences.
@@ 5.2. Text Quality
In section 2, we demonstrated that an essential characteristic of text watermarking technology is its low-impact on text quality. This means that the quality scores of texts, with or without watermarks, should be similar under the text quality evaluation function $\mathcal{R}$ as described in Equation 2. This section primarily introduces potential forms of the text quality evaluation function $\mathcal{R}$. Current text watermarking research predominantly assesses text quality using methods like perplexity values (Yang et al., 2023; Kirchenbauer et al., 2023a; Hu et al., 2023; Wu et al., 2023; Wang et al., 2023b; Zhao et al., 2023a; Liu et al., 2023c, b), semantic score (Munyer and Zhong, 2023; Yoo et al., 2023b; Yang et al., 2022, 2023; Abdelnabi and Fritz, 2021; Zhang et al., 2023a), performance evaluations for specific tasks (Topkara et al., 2006a; Zhang et al., 2023a; Hu et al., 2023; Wu et al., 2023; Yang et al., 2023; Lee et al., 2023; Sun et al., 2023), or text diversity (Kirchenbauer et al., 2023c).
@@ 5.2.1. Perplexity
The perplexity of watermarked text is generally higher than that of non-watermarked text, indicating a slight decrease in text quality. For instance, through lexical substitution, Yang et al. (2023) have generated watermarked texts whose PPL, in both Chinese and English, is marginally higher than the original texts. This same result also applies to the watermark algorithm for LLMs, such as the KGW watermark method by (Kirchenbauer et al., 2023a). However, it’s noted that the KGW algorithm does not affect the PPL of texts with very high or very low entropy. And the impact on PPL is associated with the proportion of ‘green tokens’ and the watermark’s strength; a stronger watermark strength correlates with lower text quality but increases detection success and robustness. This demonstrates a trade-off between watermark effectiveness and its impact on text quality, a conclusion applicable to other algorithms (Wang et al., 2023b; Zhao et al., 2023a; Liu et al., 2023c, b).
@@ 5.2.3. Task-specified Evaluation
Although the assessment method based on semantic scores could effectively evaluate whether adding watermarks alters the text semantics, its impact on real-world applications remains unclear. Consequently, many studies are now focusing on exploring the effects of watermarking algorithms on specific downstream tasks to assess their impact on text quality. These specific downstream tasks include machine translation (Topkara et al., 2006a; Hu et al., 2023; Wu et al., 2023; Zhao et al., 2023b), sentiment classification (Yang et al., 2023), knowledge understanding (Tu et al., 2023), code generation (Lee et al., 2023; Sun et al., 2023; Tu et al., 2023), text summarization (Hu et al., 2023; Wu et al., 2023; Tu et al., 2023), story generation (Zhao et al., 2023b), question answering (Tu et al., 2023) and instruction following (Tu et al., 2023).
@@
In the work Kirchenbauer et al. (2023c), it was discovered that for the KGW watermarking algorithm (Kirchenbauer et al., 2023a), the context width (i.e., the number of tokens used on the left to hash and generate the green list) has the most significant impact on text diversity. A larger context width enhances the diversity of the text, but this comes at the cost of reduced robustness of the watermark to text modifications. Furthermore, with a larger context width, as the watermark strength increases, so does the diversity. Conversely, with a smaller context width, an increase in watermark strength leads to a decrease in diversity. Currently, only a few studies on watermarking for Large Language Models have assessed its impact on text diversity. We suggest that future work could focus more on evaluating the effect of watermarking on diversity.
@@ 5.3. Robustness
Current assumptions for watermark removal attacks are based on black-box access to the watermarking algorithm, meaning the method of watermark generation is unknown and there is no access to the watermark detector during the attack. This is primarily because, under white-box access, potent watermark removal attack algorithms can be easily developed to remove most watermarks, as mentioned by (Kirchenbauer et al., 2023c) in their anti-watermark schema. This represents a limitation of all current watermarking algorithms. Although it is possible that watermarks may not withstand watermark removal attacks under white-box access, exploring these possibilities or designing robust watermark algorithms under white-box settings remains a direction for future research.
@@
There are some designs to further increase algorithm robustness against rewrite attack as well. 1) Decrease the token-level dependency. The dependency of watermark algorithms on specific contexts can compromise their effectiveness when the original text is rewritten, as the original context may not remain intact. Take the watermarking during logits generation methods as example, if the partitioning of a token’s red-green list relies on $m$ adjacent words, the detection rate significantly decreases after an attack when $m$ is not very small. Hence, watermark algorithms should reduce reliance on neighboring words. This can be achieved by decreasing the window size $m$ (Kirchenbauer et al., 2023a), or by using hashing schemes that only involve a subset of words within the window (Kirchenbauer et al., 2023c). Some studies have even reduced the window size $m$ to zero (Zhao et al., 2023a). 2) Replace with semantic-level dependency. A better approach is for the watermark algorithm to refer to the general semantics around the watermarked token (Liu et al., 2023c; Yoo et al., 2023b) instead of the exactly identical tokens or words. That is to say, for a group of altered but semantically similar contexts, the watermark generator and detector would produce the same result. Unless the semantics is drastically altered, the watermark detection could maintain a higher accuracy against rewrite attacks. One can even train the watermark generator to produce closer results on semantically close texts ; 3) Increase embedding space for each message. The paraphrasing attack would change the appearance of a watermarked text fragment and disable accurate detection in that fragment. If the message is only embedded in one short text piece, the message would be easily erased if the re-writer considered that part redundant. Thus, increasing the number or length of text fragments that are embedded with the same message would intuitively increase the success extraction rate (Munyer and Zhong, 2023; Wang et al., 2023b).
@@
The copy-paste attack primarily diminishes the detectability of watermarks in text by diluting their concentration. The efficacy of this attack is significantly influenced by the proportion of watermarked text. For instance, when the watermarked text constitutes a mere 10% of the total, the attack’s impact tends to surpass that of most rewriting attacks (Kirchenbauer et al., 2023c). However, as the share of watermarked text increases to, say, 25%, its effectiveness in undermining certain watermarking algorithms (Kirchenbauer et al., 2023a) becomes comparable to that of some rewriting attacks. Similar to rewriting attacks, lengthening the text can enhance the reliability of watermark detection, particularly in the context of copy-paste attacks.
However, copy-paste attacks may be identified by certain specific watermark detection methods. For example, Kirchenbauer et al. (2023c) mentioned a windowed test to calculate the level of watermarking in a region of the text instead of the whole text. The idea is to detect watermarked text that is inserted into an existing text. This should be effective for copy-paste attack.
@@ 5.4. Unforgeability
In section 5.3, we discussed watermark removal attacks that assume the attacker is unaware of the watermark’s generation method (black-box setting). However, in real-world scenarios, attackers may attempt to obtain or decipher the watermark’s generation method. Once an attacker acquires this method, they can easily fabricate or remove existing watermarks, as mentioned by (Kirchenbauer et al., 2023c) in their anti-watermark methods. Therefore, a watermark algorithm must possess substantial unforgeability, meaning it should be exceedingly difficult for attackers to discern the method of watermark generation. This might imply that the algorithm itself needs to be highly complex and secure, or involve mathematical problems that are challenging to crack.

### v3

@@ 4.3. Watermarking during Token Sampling
The previous section primarily focused on incorporating watermarks during the logits generation phase for Large Language Models. However, even after embedding watermarks, it is necessary to sample the next token from the watermarked logits. Consequently, the effectiveness of the watermark might be influenced by different sampling methods; for instance, beam search typically allows for a higher watermark intensity. In this section, we will introduce a technique of watermarking during token sampling, which does not alter the logits produced by the LLM. The primary advantage of this method is that the generated text is usually unbiased, thus minimally affecting text quality and serving as the first line of defense against watermark forgery.
@@
To address these issues, Kuditipudi et al. (2023) proposed the use of a pseudo-random number sequence significantly longer than the text, randomly selecting a starting position from the sequence for each watermark insertion to introduce randomness. Additionally, during watermark detection, they incorporate a soft notion of edit distance (i.e., Levenshtein distance) into the computation of the alignment between text and the pseudo-random number sequence. This approach significantly enhances the robustness of the watermarking algorithm against watermark removal attacks. Apart from intervening in the sampling process of each token one by one, Hou et al. (2023) suggested incorporating watermarks during sentence-level sampling. The algorithm initially partitions the semantic embedding space into a watermarked region and a non-watermarked region, and then performs sentence-level rejection sampling until the sampled sentence falls within the watermarked region. Since the partition principles are based on sentence-level semantics, this approach significantly enhances robustness against watermark removal attacks such as paraphrasing. Currently, there is limited research on watermarking during token sampling, suggesting substantial potential for growth in this field. Moreover, the practical effectiveness and robustness of this method may require further validation through more experiments and real-world applications.
@@ 5. Evaluation Metrics for Text Watermarking
In sections 3 and 4, we have provided a comprehensive introduction to the existing text watermarking methods. Meanwhile, for a text watermarking algorithm, it is crucial to conduct a comprehensive evaluation from various perspectives. In this section, we will introduce how to evaluate a watermarking algorithm from four perspectives: success rate, text quality, robustness and unforgeability. Among these, success rate (section 5.1) refers to the ability to correctly detect watermarked texts. Text quality (section 5.2) assesses whether the quality of watermarked text is degraded compared to non-watermarked text. Robustness (section 5.3) examines whether the watermarked text can still be detected after watermark removal attacks. Finally, unforgeability (section 5.4) evaluates the difficulty for third parties to forge the watermark. Furthermore, Table 1 outlines the alignment between various text watermarking algorithms discussed in the survey and the evaluation perspectives they contribute to.
Table 1. Relationships between text watermarking algorithms covered in the survey and the evaluation metrics, featuring the individual objectives each text watermarking algorithm aims to achieve. $\blacktriangle$ stands for basic objectives, $\bullet$ stands for primary objectives, and $\circ$ stands for secondary objectives.
Text Watermarking Algorithms Objectives
Watermarked
Object Category Method Success Rate Text Quality Robustness Unforgeability
Detection Accuracy Payload
Existing Text
(section˜3) Format-based
(section˜3.1) Line/Word-Shift Coding (Brassil et al., 1995) $\blacktriangle$ $\bullet$
UniSpach (Por et al., 2012) $\bullet$ $\bullet$
Unicode Homoglyph Substitution (Rizzo et al., 2016) $\bullet$ $\bullet$
EasyMark (Sato et al., 2023) $\bullet$ $\bullet$
Lexical-based
(section˜3.2) Equimark (Topkara et al., 2006b) $\blacktriangle$ $\bullet$ $\circ$
DeepTextMark (Munyer and Zhong, 2023) $\bullet$ $\bullet$
Context-aware Lexical Substitution (Yang et al., 2022) $\circ$ $\bullet$ $\circ$
Binary-encoding Lexical Substitution (Yang et al., 2023) $\circ$ $\circ$
Robust Multi-bit Watermark (Yoo et al., 2023b) $\circ$ $\circ$ $\bullet$
Syntatic-based
(section˜3.3) NLW (Atallah et al., 2001) $\blacktriangle$ $\bullet$
WANE (Topkara et al., 2006a) $\bullet$ $\circ$ $\bullet$
MA-NLW (Meral et al., 2009) $\bullet$ $\circ$ $\circ$
Generation-based
( section˜3.4) AWT (Abdelnabi and Fritz, 2021) $\blacktriangle$ $\bullet$ $\bullet$
REMARK-LLM (Zhang et al., 2023a) $\bullet$ $\circ$ $\circ$
LLMs
(section˜4) Training Time
(section˜4.1) Dataset Trigger (Liu et al., 2023a) $\blacktriangle$ $\bullet$
Clean-Label Backdoor Watermark (Tang et al., 2023) $\bullet$
Coprotector (Sun et al., 2022) $\bullet$
CodeMark (Sun et al., 2023) $\bullet$
Logits Generation
(section˜4.2) KGW (Kirchenbauer et al., 2023a) $\blacktriangle$ $\bullet$
SWEET (Lee et al., 2023) $\bullet$
Unbiased Watermark (Hu et al., 2023) $\bullet$ $\circ$ $\bullet$
DiPmark (Wu et al., 2023) $\bullet$ $\circ$ $\bullet$
COLOR (Yoo et al., 2023a) $\bullet$ $\circ$
CTWL (Wang et al., 2023b) $\bullet$ $\bullet$ $\circ$ $\bullet$
ThreeBrick (Fernandez et al., 2023) $\circ$
Unigram Watermark (Zhao et al., 2023a) $\bullet$
KGW-reliability (Kirchenbauer et al., 2023b) $\bullet$
NS-Watermark (Takezawa et al., 2023) $\bullet$ $\circ$
Semantic Invariant Robust Watermark (Liu et al., 2023c) $\circ$ $\bullet$ $\bullet$
Unforgeable Watermark (Liu et al., 2023b) $\circ$ $\circ$ $\bullet$
Publicly Detectable Watermark (Fairoze et al., 2023) $\circ$ $\circ$ $\bullet$
Semantic-based Robust Watermark (Ren et al., 2023) $\circ$ $\bullet$
Token Sampling
(section˜4.3) Undetectable Watermark (Christ et al., 2023) $\blacktriangle$ $\bullet$ $\circ$ $\bullet$
Robust Distortion-free Watermark (Kuditipudi et al., 2023) $\bullet$ $\bullet$ $\bullet$
SemStamp (Hou et al., 2023) $\circ$ $\bullet$
@@ 5.1.1. Zero-bit Watermark
In zero-bit watermarking, the watermarking algorithm can only determine whether a text contains a watermark, without the ability to extract additional information from the watermarked text. Given that the detection of a zero-bit watermark inherently represents a binary classification problem — discerning whether a text is watermarked or not — the evaluation of its success rate typically employs classification metrics, such as accuracy and F1 score. In most works (Kirchenbauer et al., 2023a; Zhao et al., 2023a; Liu et al., 2023b, c), since the proportion of watermarked and non-watermarked texts in the test dataset is usually equal, the values of accuracy and F1 score tend to be quite similar. Despite this, the F1 score is often more frequently utilized, accompanied by additional metrics including false positive and false negative rate (Kirchenbauer et al., 2023a; Zhao et al., 2023a; Liu et al., 2023b). Here, false positives denote the erroneous classification of non-watermarked texts as watermarked, whereas false negatives refer to the incorrect classification of watermarked texts as non-watermarked. Generally, false positives are more important since misidentifying human-generated texts as watermarked can lead to more adverse consequences.
@@ 5.2. Text Quality
In section 2, we demonstrated that an essential characteristic of text watermarking technology is its low-impact on text quality. This means that the quality scores of texts, with or without watermarks, should be similar under the text quality evaluation function $\mathcal{R}$ as described in Equation 2. This section primarily introduces potential forms of the text quality evaluation function $\mathcal{R}$. Current text watermarking research predominantly assesses text quality using methods like perplexity values (Yang et al., 2023; Kirchenbauer et al., 2023a; Hu et al., 2023; Wu et al., 2023; Wang et al., 2023b; Zhao et al., 2023a; Liu et al., 2023c, b), semantic score (Munyer and Zhong, 2023; Yoo et al., 2023b; Yang et al., 2022, 2023; Abdelnabi and Fritz, 2021; Zhang et al., 2023a), performance evaluations for specific tasks (Topkara et al., 2006a; Zhang et al., 2023a; Hu et al., 2023; Wu et al., 2023; Yang et al., 2023; Lee et al., 2023; Sun et al., 2023), or text diversity (Kirchenbauer et al., 2023b).
@@ 5.2.1. Perplexity
The perplexity of watermarked text is generally higher than that of non-watermarked text, indicating a slight decrease in text quality. Generally, the impact of watermarking algorithms on text quality is correlated with the strength of the watermark. The higher the watermark strength, the more evident the decline in text quality. For instance, in the KGW algorithm (Kirchenbauer et al., 2023a), a higher value of $\delta$ results in greater impact on text quality. Takezawa et al. (2023) suggest that for longer texts, a weaker watermark strength can be employed to minimize the effect on text quality while maintaining the effectiveness of the watermark.
@@ 5.2.3. Task-specified Evaluation
Although the assessment method based on semantic scores could effectively evaluate whether adding watermarks alters the text semantics, its impact on real-world applications remains unclear. Consequently, many studies are now focusing on exploring the effects of watermarking algorithms on specific downstream tasks to assess their impact on text quality. These specific downstream tasks include machine translation (Topkara et al., 2006a; Hu et al., 2023; Wu et al., 2023; Zhao et al., 2023b), sentiment classification (Yang et al., 2023), knowledge understanding (Tu et al., 2023), code generation (Lee et al., 2023; Sun et al., 2023; Tu et al., 2023), text summarization (Hu et al., 2023; Wu et al., 2023; Tu et al., 2023), story generation (Zhao et al., 2023b), question answering (Tu et al., 2023) and instruction following (Tu et al., 2023), as illustrated in Figure 6.
@@
Figure 6. Specific downstream tasks used to evaluate the impact of text watermarking algorithms on text quality.
@@
In the work Kirchenbauer et al. (2023b), it was discovered that for the KGW watermarking algorithm (Kirchenbauer et al., 2023a), the context width (i.e., the number of tokens used on the left to hash and generate the green list) has the most significant impact on text diversity. A larger context width enhances the diversity of the text, but this comes at the cost of reduced robustness of the watermark to text modifications. Furthermore, with a larger context width, as the watermark strength increases, so does the diversity. Conversely, with a smaller context width, an increase in watermark strength leads to a decrease in diversity. Currently, only a few studies on watermarking for Large Language Models have assessed its impact on text diversity. We suggest that future work could focus more on evaluating the effect of watermarking on diversity.
@@ 5.3. Robustness
Current assumptions for watermark removal attacks are based on black-box access to the watermarking algorithm, meaning the method of watermark generation is unknown and there is no access to the watermark detector during the attack. This is primarily because, under white-box access, potent watermark removal attack algorithms can be easily developed to remove most watermarks, as mentioned by (Kirchenbauer et al., 2023b) in their anti-watermark schema. This represents a limitation of all current watermarking algorithms. Although it is possible that watermarks may not withstand watermark removal attacks under white-box access, exploring these possibilities or designing robust watermark algorithms under white-box settings remains a direction for future research.
@@
There are some designs to further increase algorithm robustness against rewrite attack as well. 1) Decrease the token-level dependency. The dependency of watermark algorithms on specific contexts can compromise their effectiveness when the original text is rewritten, as the original context may not remain intact. Take the watermarking during logits generation methods as example, if the partitioning of a token’s red-green list relies on $m$ adjacent words, the detection rate significantly decreases after an attack when $m$ is not very small. Hence, watermark algorithms should reduce reliance on neighboring words. This can be achieved by decreasing the window size $m$ (Kirchenbauer et al., 2023a), or by using hashing schemes that only involve a subset of words within the window (Kirchenbauer et al., 2023b). Some studies have even reduced the window size $m$ to zero (Zhao et al., 2023a). 2) Replace with semantic-level dependency. A better approach is for the watermark algorithm to refer to the general semantics around the watermarked token (Liu et al., 2023c; Yoo et al., 2023b) instead of the exactly identical tokens or words. That is to say, for a group of altered but semantically similar contexts, the watermark generator and detector would produce the same result. Unless the semantics is drastically altered, the watermark detection could maintain a higher accuracy against rewrite attacks. One can even train the watermark generator to produce closer results on semantically close texts ; 3) Increase embedding space for each message. The paraphrasing attack would change the appearance of a watermarked text fragment and disable accurate detection in that fragment. If the message is only embedded in one short text piece, the message would be easily erased if the re-writer considered that part redundant. Thus, increasing the number or length of text fragments that are embedded with the same message would intuitively increase the success extraction rate (Munyer and Zhong, 2023; Wang et al., 2023b).
@@
The copy-paste attack primarily diminishes the detectability of watermarks in text by diluting their concentration. The efficacy of this attack is significantly influenced by the proportion of watermarked text. For instance, when the watermarked text constitutes a mere 10% of the total, the attack’s impact tends to surpass that of most rewriting attacks (Kirchenbauer et al., 2023b). However, as the share of watermarked text increases to, say, 25%, its effectiveness in undermining certain watermarking algorithms (Kirchenbauer et al., 2023a) becomes comparable to that of some rewriting attacks. Similar to rewriting attacks, lengthening the text can enhance the reliability of watermark detection, particularly in the context of copy-paste attacks.
However, copy-paste attacks may be identified by certain specific watermark detection methods. For example, Kirchenbauer et al. (2023b) mentioned a windowed test to calculate the level of watermarking in a region of the text instead of the whole text. The idea is to detect watermarked text that is inserted into an existing text. This should be effective for copy-paste attack.
@@ 5.4. Unforgeability
In section 5.3, we discussed watermark removal attacks that assume the attacker is unaware of the watermark’s generation method (black-box setting). However, in real-world scenarios, attackers may attempt to obtain or decipher the watermark’s generation method. Once an attacker acquires this method, they can easily fabricate or remove existing watermarks, as mentioned by (Kirchenbauer et al., 2023b) in their anti-watermark methods. Therefore, a watermark algorithm must possess substantial unforgeability, meaning it should be exceedingly difficult for attackers to discern the method of watermark generation. This might imply that the algorithm itself needs to be highly complex and secure, or involve mathematical problems that are challenging to crack.
@@ 6.3. Fake News Detection
7. Challenges and Future Directions
Although detailed introductions to the methods, evaluations, and application scenarios of text watermarking have been provided in previous sections, numerous challenges remain in this field. These include balancing across different evaluation perspectives, adapting text watermarking for more challenging scenarios, developing more comprehensive benchmarks, and broadening the application of text watermarking. These challenges will be discussed in detail below.
7.1. Balancing Across Different Evaluation Perspectives
In section 5, we explore various perspectives for evaluating text watermarking algorithms. However, these perspectives often present inherent contradictions, making it extremely challenging for a text watermarking algorithm to excel in all evaluation perspectives simultaneously. For instance, achieving a favorable balance among success rate, text quality, and robustness at a high payload is difficult. In this section, we will first analyze why these perspectives are mutually contradictory, and then discuss potential strategies for achieving a better balance in future work.
7.1.1. Why are the Different Perspectives Conflicting?
The fundamental reason for contradictions among different perspectives lies in the limited suitable text space for text watermarking, usually determined by the text quality requirements. Specifically, according to Equation 2, the score difference between watermarked and non-watermarked texts under the quality evaluation function $\mathcal{R}$ should be less than a threshold $\beta$. However, the number of texts meeting this criterion is limited, denoted as $|t_{\beta}|$. Since the minimal impact on text quality is a crucial feature of text watermarking algorithms, there is an upper limit for $|t_{\beta}|$ for all watermarking algorithms. Given the watermark text space of $|t_{\beta}|$, we can further analyze the conflicts between different evaluation perspectives.
We begin by introducing conflicts among success rate, text quality, and robustness at a high payload. Payload and robustness involve different strategies for partitioning this limited text space. If the space is divided to encode more watermark messages (i.e., a larger payload), minor modifications to any watermarked text are more likely to result in detection as another watermark message, thus reducing robustness. Conversely, reducing the number of watermark messages encoded (i.e., lower payload) decreases the likelihood that modifications to watermarked text result in other watermark information, thereby increasing robustness. Hence, the conflict between payload and robustness is evident. Simultaneously, as text quality requirements increase, the size of the text space $|t_{\beta}|$ diminishes, potentially leading to a decrease in both payload and robustness, making the conflict between text quality and these two metrics obvious.
Lastly, we analyze why unforgeability and other evaluation perspectives might conflict. Typically, enhancing the complexity of watermarking algorithms can improve their unforgeability. However, incorporating additional complex modules into the algorithm often introduces greater robustness risks. Moreover, in publicly detectable scenarios, algorithms seeking to enhance unforgeability often conceal their detection methods using some watermark text space (Fairoze et al., 2023; Liu et al., 2023b), thereby conflicting with other evaluation approaches.
7.1.2. Future Directions
The current text watermarking algorithms primarily focus on balancing robustness and the impact on text quality. However, these efforts often do not simultaneously address payload and unforgeability, which should be the main focus of future work.
The key to balancing payload, robustness, and text quality lies primarily in devising a more effective strategy for partitioning the watermark text space. This may require additional designs to counter potential watermark removal attacks, dividing the watermark space into different watermark messages, ensuring that transitioning between different watermark messages necessitates a sufficient number of watermark removal attack operations. Secondly, from the perspective of the payload, it is feasible to draw inspiration from the concepts of error-correction codes, such as utilizing Hamming codes (Kumar and Umashankar, 2007), to enhance the probability of recovering the original watermark information from partially modified text. These methods may effectively enhance payload and robustness while maintaining a consistent impact on text quality.
To enhance the unforgeability of text watermarks, it is generally necessary to utilize expertise from fields such as cryptography, information theory, and machine learning. This involves increasing the complexity of watermarking algorithms to improve their resistance to forgery. Although current methods have made some progress, their more intricate designs still introduce additional non-robust factors. Furthermore, these methods have not been extended to scenarios with larger payloads.
7.2. Adapting Text Watermarking for More Challenging Scenarios
Current watermark algorithms often achieve satisfactory results in relatively simple contexts. However, they require further enhancement and adaptation when confronted with more challenging scenarios, such as low-entropy scenarios and publicly detectable scenarios.
In low-entropy situations like coding, legal, and medical texts, diversity and complexity are typically reduced. These types of texts typically adhere to strict formatting (grammatical) requirements, so embedding watermarks without affecting these requirements is challenging. A more in-depth explanation is that for low-entropy text, the upper limit of the watermark text space is lower, thereby making it harder to find suitable watermark text. Future methods may need a stronger understanding of their formatting or grammatical requirements, and thereby designing semantically invariant format transformations to expand the available watermark text space.
In scenarios where watermarks are publicly detectable, both the presence of the watermark and its detection mechanism are openly visible. This poses greater challenges for the design of watermark algorithms. Firstly, the algorithm must be sufficiently complex and unpredictable to ensure that, even in a public setting, attackers struggle to effectively corrupt or alter the watermark information. Secondly, the design must ensure that the watermark’s generation method cannot be inferred even through its detector. This primarily involves considerations of unforgeability, but in publicly detectable scenarios, there are also heightened demands for robustness and text quality. Future methods will need to integrate considerations of security and practicality, potentially involving more intricate encryption and machine learning technologies.
7.3. Developing More Comprehensive Benchmarks
Current research in text watermarking benchmark primarily focuses on text quality, with limited benchmarks for other critical metrics like high success rate with large payload, robustness, and unforgeability. Therefore, developing a more comprehensive benchmarking system is a crucial direction for future research. Constructing such benchmarks requires significant effort, taking into account various application scenarios, attack methods, and characteristics of different watermarking algorithms. It also necessitates establishing a fair, transparent, and user-friendly evaluation process, allowing researchers to test and compare algorithms under unified standards. This benchmarking system will not only advance academic research into text watermarking algorithms but also aid the industry in better understanding and applying these technologies.
7.4. Broadening the Application of Text Watermarking
Although text watermarking technology has demonstrated its practicality in multiple domains, its wider application necessitates further efforts. This encompasses not only advancements in watermarking techniques but also factors beyond the technical realm. In this section, we will explore the challenges faced in expanding the applications of text watermarking technology, particularly from the perspectives of large language model providers, public trust, and transparency.
7.4.1. Limited Engagement of Large Language Model Providers
As an increasing amount of text is generated directly by large language models, it is crucial for providers of these LLMs to integrate text watermarking functionality into their services to promote the use of text watermarks. However, the current level of engagement from these providers in text watermarking technology remains insufficient, influenced by various technical and non-technical factors.
Firstly, current text watermarking algorithms cannot guarantee no reduction in text quality, which may impact the service quality of model providers. Technically, this demands future text watermarking algorithms to consider the impact on text quality more thoroughly. Additionally, the direct benefit for LLM providers from text watermarking technology lies in the protection of the LLMs’ copyrights. More focus on research in this area is needed to encourage providers to participate more actively in promoting text watermarking technology.
Moreover, non-technical factors also affect the engagement level of large model providers. For instance, the role of governments and regulatory bodies is significant. There needs to be a discussion on how governments should use legal restrictions or incentives to encourage model providers to adopt watermarking technology. This could involve setting relevant standards and norms, as well as providing economic incentives to encourage the adoption of these technologies.
7.4.2. Lack of Public Trust
Public trust and transparency in text watermarking technology are key factors in promoting its widespread application. Only when the public trusts the text watermarking algorithms and believes in the accuracy of their detection results can they be effectively utilized in practical applications. To enhance public trust, it is necessary to ensure the transparency and reliability of watermarking technology.
A fundamental step towards this goal involves the comprehensive disclosure of the text watermarking detection algorithms. Making these details accessible allows users to grasp and assess the principles and accuracy of the algorithms. Such transparency not only cultivates user trust but also spurs further academic and industrial advancements.
Moreover, involving independent third-party platforms for detection and verification can strengthen trust. These platforms offer unbiased evaluations, alleviating conflict of interest concerns. Furthermore, government and regulatory guidelines can ensure the technology’s fairness and transparency, further boosting public confidence.